Which pair has equally likely outcomes?
a. drawing a red 5 out of a standard 52 card deck given it is not a face card or ace.
b. rolling a total of 10 on a two fair six sided dice.
c. rolling a total of 3 on a two fair six sided dice.
d. rolling a total of 7 on a two fair six sided dice.
e. drawing a four out of a standard 52 card deck given it is not a face card or an ace.
** I have checked all probabilities and can't seem to find two equal ones. I have to be doing something wrong, please help AND thank you in advance for your help!!!

Answer by Theo(13342)

I believe your solution is as follows:

If you draw a red 5 out of a standard 52 card deck given it is not a face card or an ace, then the probability of success would be 2/32 = 1/18.

This is because you are given that the card drawn is not a face card or an ace.

There are 4 face cards plus an ace for each suit in the deck.

That's 5 cards per suit * 4 for a total of 20 cards that can be effectively removed from the deck because you are given that the card drawn will not be a face card or an ace.

Your universe of possible selections is now 32 rather than 52 because you are given that the card will not be a face card or an ace.

The matching probability would be drawing a total of 3 out of 2 dice.

The probability that can happen is 2 / 36 = 1 / 18 which is the same as drawing the red card out of a deck without the face cards or the ace in them.